# Capillary sorting of fiber suspensions by dip coating

### v1

@@ Abstract
A potential solution to filter anisotropic particles, specifically fibers, is capillary filtration using a dip coating process. The principle is that a substrate is submersed within a suspension bath, and based on the withdrawal velocity of the substrate and the properties of the interstitial fluid, only particles below a certain threshold size will be entrained in the liquid film coating the substrate after removal Gans et al. (2019); Palma and Lhuissier (2019); Shovon et al. (2022); Copeland et al. (2023). Previous works have demonstrated the efficacy of sorting spherical particles using dip coating with a planar substrate Sauret et al. (2019); Dincau et al. (2019) and cylindrical substrates Dincau et al. (2020); Khalil and Khoda (2022).
@@
Methods. Schematics of the experimental setup are shown in Figs. 1(a)-(b). The suspensions are made of non-Brownian nylon fibers of diameter $200\,\mu{\rm m}\leq d\leq 400\,\mu{\rm m}$ and length $1.5\,{\rm mm}\leq L\leq 4.5\,{\rm mm}$ (standard deviation of $0.25\,{\rm mm}$) dispersed in Fluorosilicone (FMS-221, Gelest), a Newtonian fluid of viscosity $\mu_{\rm f}=109\,{\rm mPa.s}$, surface tension $\gamma=21\,\mathrm{mN.m^{-1}}$ and density $\rho_{\rm f}=1160\,\mathrm{kg.m^{-3}}$. The fibers used are cut to the desired length by mechanical chopping. Therefore, the edges of the fibers are not perfectly cylindrical, but, for instance, a little bit flattened irregularly. Nevertheless, we did not observe any effects of the edge geometry on the sorting process. The density of the liquid is close enough to the density of the fibers ($\rho\simeq 1150\,\mathrm{kg.m^{-3}}$) so that buoyancy effects can be neglected over the timescale of the experiments. The volume fraction of fibers used, $\phi=V_{\rm fib}/V_{\rm tot}=0.5\%$, ensures that we remain in the dilute regime Butler and Snook (2018). The suspensions are initially mixed overnight using magnetic stirrer bars and mixed for 5 minutes between trials. Using optical absorbance, we ensured that our experimental protocol led to fibers initially homogeneously dispersed in the suspension bath.
Figure 1: Schematic of the dip coating process with a suspension of fibers of diameter $d$ and two different lengths $L_{1}$ and $L_{2}$. (a) Side and (b) top view. Note that the size of the fibers, the substrate, and the meniscus are not to scale.
We perform the experiments with two substrate geometries: a planar substrate that consists of a glass plate (7mm thick, 50mm wide, 70 mm length) and a cylindrical substrate setup consisting of an array of three rods (each of radius $R_{\rm s}=1.2\,{\rm mm}$), withdrawn simultaneously from the suspension bath. The substrates are withdrawn from the suspension bath using a linear translation stage (Thorlabs NRT150 linear stepper motor) at a velocity ranging from $2.5\,{\rm mm.s^{-1}}$ to $20\,{\rm mm.s^{-1}}$ over a distance of $30\,{\rm mm}$. In both cases, between trials, the substrates were cleaned using isopropyl alcohol, rinsed with Deionized (DI) water, and dried using compressed air.
@@
Discussion. The fibers used here are of length $L\gg h$, so that they will need some amount of reorientation to fit within the meniscus as viscous forces pull them upward or else be subjected to surface forces, which will push them back downward. In addition, a key difference between planar and cylindrical substrates is the curvature in the horizontal direction (around the cylindrical substrates). Past experiments have shown that planar substrates allow for most fibers to have their axis becomes along the meniscus near the entrainment threshold Jeong et al. (2023), thus allowing the coupling between capillary effects and the frictional force. Conversely, a cylindrical substrate has a curved film, which means that long fibers ($L^{*}\gtrsim 1$) must reorient such that their main axis becomes parallel to the axis of the substrate to become entrained. This means that for a cylindrical substrate, longer fibers must take a path of higher resistance to entrainment, fully rotating to the preferred orientation, and without an alternative lower resistance path like the planar substrate, longer fibers will be consistently repelled. As a result, the threshold capillary numbers for longer fibers ($L^{*}\gtrsim 1$) increases with $L^{*}$. The consequence of this difference is that attempting to sort by length with planar substrates is less efficient, as illustrated in Fig. 5. However, with cylindrical substrates, a better sorting efficiency can be achieved, as reported in Fig. 5 where when attempting to sort fibers of $200\,\mu{\rm m}$ diameter and length between $1.5\,{\rm mm}$ and $3.5\,{\rm mm}$, a planar substrate (where $L^{*}$ = 0) resulted in a peak efficiency of 36%, while a $R_{\rm s}=1.2\,{\rm mm}$ radius cylindrical substrate (where $L^{*}$ = 0.6 for the $1.5\,{\rm mm}$ fibers and $L^{*}$ = 1.4 for the $3.5\,{\rm mm}$ fibers) resulted in a peak efficiency of 100%. We should emphasize that the difference in operating capillary numbers between the two substrates is due to the different evolution of the film thickness with ${\rm Ca}$, as shown in Eqs. (1) and (2) for the flat and cylindrical substrates, respectively.
@@
Even when a fiber suspension has a large enough disparity in length to ideally sort perfectly, volume fraction effects can lead to a decrease in efficiency due to the entrainment of clusters of particles Sauret et al. (2019). Indeed, when the local volume fraction at the meniscus is too high, collective effects may appear, and the desired length fibers will form clusters that can become tangled with the undesired length and pass through the meniscus as one large aggregate Jeong et al. (2022). Reducing the average volume fraction of the entire suspension will help to minimize this effect. Ideally, the volume fraction would be as low as possible, however, very low volume fractions lead to a much smaller average number of particles entrained per trial, which may be undesirable to create a rapid filtering process. As such, the volume fraction chosen should provide an acceptable balance between filtering efficiency and process duration to meet their individual requirements.

### v2

@@ Abstract
A potential solution to filter anisotropic particles, specifically fibers, is capillary filtration using a dip coating process. The principle is that a substrate is submersed within a suspension bath, and based on the withdrawal velocity of the substrate and the properties of the interstitial fluid, only particles below a certain threshold size will be entrained in the liquid film coating the substrate after removal Gans et al. (2019); Palma and Lhuissier (2019); Shovon et al. (2022); Copeland et al. (2023). Previous works have demonstrated the efficacy of sorting spherical particles using dip coating with a planar substrate Sauret et al. (2019); Dincau et al. (2019) and cylindrical substrates Dincau et al. (2020); Khalil and Khoda (2022). These sorting methods rely on the importance of the size of the particle compared to the size of the capillary film. Such interplay between liquid/air interfaces and solid particles have also been considered for liquid thread Furbank and Morris (2004); Bonnoit et al. (2012); Château et al. (2018); Thiévenaz et al. (2021); Thiévenaz and Sauret (2022) and films Raux et al. (2020); Jeong et al. (2022a).
@@
Methods. Schematics of the experimental setup are shown in Figs. 1(a)-(b). The suspensions are made of non-Brownian nylon fibers of diameter $200\,\mu{\rm m}\leq d\leq 400\,\mu{\rm m}$ and length $1.5\,{\rm mm}\leq L\leq 4.5\,{\rm mm}$ dispersed in Fluorosilicone (FMS-221, Gelest), a Newtonian fluid of viscosity $\mu_{\rm f}=109\,{\rm mPa.s}$, surface tension $\gamma=21\,\mathrm{mN.m^{-1}}$ and density $\rho_{\rm f}=1160\,\mathrm{kg.m^{-3}}$. The fibers used are cut to the desired length by mechanical chopping (WC302 Automatic Wire Cutter, The Eraser Company), resulting in a well-controlled average length $L$ and a standard deviation on the length of $0.25\,{\rm mm}$. Due to the mechanical chopping, the edges of the fibers are not perfectly cylindrical, but, for instance, a little bit flattened irregularly. Nevertheless, we did not observe any effects of the edge geometry on the sorting process. In our experiments, no deformation of the fibers was observed for any aspect ratio considered. The density of the liquid is close enough to the density of the fibers ($\rho\simeq 1150\,\mathrm{kg.m^{-3}}$) so that buoyancy effects can be neglected over the timescale of the experiments. The volume fraction of fibers used, $\phi=V_{\rm fib}/V_{\rm tot}=0.5\%$, ensures that we remain in the dilute regime Butler and Snook (2018). The suspensions are initially mixed overnight using magnetic stirrer bars and mixed for 5 minutes between trials. Using optical absorbance, we ensured that our experimental protocol led to fibers initially homogeneously dispersed in the suspension bath.
Figure 1: (a) Side and (b) top view schematics of the dip coating process with a suspension of fibers of diameter $d$ and two different lengths $L_{1}$ and $L_{2}$. Note that the size of the fibers, the substrate, and the meniscus are not to scale. (c)-(d) Examples of entrainment of fibers on (c) a flat substrate where no preferential orientation is observed and (d) on a rod where the fibers entrained are of length comparable to the substrate diameter and thus mainly align with the axis of the rod.
We perform the experiments with two substrate geometries: a planar substrate that consists of a glass plate (7mm thick, 50mm wide, 70 mm length; see figure 1(c)) and a cylindrical substrate setup consisting of an array of three rods (each of radius $R_{\rm s}=1.2\,{\rm mm}$; see figure 1(d)), withdrawn simultaneously from the suspension bath. The substrates are withdrawn from the suspension bath using a linear translation stage (Thorlabs NRT150 linear stepper motor) at a velocity ranging from $2.5\,{\rm mm.s^{-1}}$ to $20\,{\rm mm.s^{-1}}$ over a distance of $30\,{\rm mm}$. In both cases, between trials, the substrates were cleaned using isopropyl alcohol, rinsed with Deionized (DI) water, and dried using compressed air.
@@
Discussion. The fibers used here are of length $L\gg h$, so that they will need some amount of reorientation to fit within the meniscus as viscous forces pull them upward or else be subjected to surface forces, which will push them back downward, as described in Ref. Jeong et al. (2023). In addition, a key difference between planar and cylindrical substrates is the curvature in the horizontal direction (around the cylindrical substrates). Past experiments have shown that planar substrates allow for most fibers to have their axis become along the meniscus near the entrainment threshold Jeong et al. (2023), thus allowing the coupling between capillary effects and the frictional force. Conversely, a cylindrical substrate has a curved film, which means that long fibers ($L^{*}\gtrsim 1$) must reorient such that their main axis becomes parallel to the axis of the substrate to become entrained. This means that for a cylindrical substrate, longer fibers must take a path of higher resistance to entrainment, fully rotating to the preferred orientation, and without an alternative lower resistance path like the planar substrate, longer fibers will be consistently repelled. As a result, the threshold capillary numbers for longer fibers ($L^{*}\gtrsim 1$) increases with $L^{*}$. The consequence of this difference is that attempting to sort by length with planar substrates is less efficient, as illustrated in Fig. 5. However, with cylindrical substrates, a better sorting efficiency can be achieved, as reported in Fig. 5 where when attempting to sort fibers of $200\,\mu{\rm m}$ diameter and length between $1.5\,{\rm mm}$ and $3.5\,{\rm mm}$, a planar substrate (where $L^{*}$ = 0) resulted in a peak efficiency of 36%, while a $R_{\rm s}=1.2\,{\rm mm}$ radius cylindrical substrate (where $L^{*}$ = 0.6 for the $1.5\,{\rm mm}$ fibers and $L^{*}$ = 1.4 for the $3.5\,{\rm mm}$ fibers) resulted in a peak efficiency of 100%. We should emphasize that the difference in operating capillary numbers between the two substrates is due to the different evolution of the film thickness with ${\rm Ca}$, as shown in Eqs. (1) and (2) for the flat and cylindrical substrates, respectively.
@@
Even when a fiber suspension has a large enough disparity in length to ideally sort perfectly, volume fraction effects can lead to a decrease in efficiency due to the entrainment of clusters of particles Sauret et al. (2019). Indeed, when the local volume fraction at the meniscus is too high, collective effects may appear, and the desired length fibers will form clusters that can become tangled with the undesired length and pass through the meniscus as one large aggregate Jeong et al. (2022b). Reducing the average volume fraction of the entire suspension will help to minimize this effect. Ideally, the volume fraction would be as low as possible, however, very low volume fractions lead to a much smaller average number of particles entrained per trial, which may be undesirable to create a rapid filtering process. As such, the volume fraction chosen should provide an acceptable balance between filtering efficiency and process duration to meet their individual requirements.